OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 + 19x + 84 A. x - 7 B. x2 + 12 C. x + 84 D. x + 7

OpenStudy (***[isuru]***):

u can factorize this in this way x2 + 19x + 84 = x2 + 12 x + 7x + 84 = (x + 7) ( x+ 12 ) now u can find ur answer :)

OpenStudy (anonymous):

THANKS.

OpenStudy (anonymous):

how about this one. Which of the binomials below is a factor of this trinomial? x2 + 6x + 8 A. x + 8 B. x2 + 4 C. x - 4 D. x + 4

OpenStudy (***[isuru]***):

this one can factorize as (x + 4)( x+ 2) = x2 + 4x + 2x + 8 = x2 + 6x + 8 :p